# Evaluating the Efficacy of Prompt-Engineered Large Multimodal Models Versus Fine-Tuned Vision Transformers in Image-Based Security Applications

### v1

@@ Abstract
The success of Large Language Models (LLMs) has led to a parallel rise in the development of Large Multimodal Models (LMMs), such as Gemini-pro, which have begun to transform a variety of applications. These sophisticated multimodal models are designed to interpret and analyze complex data, integrating both textual and visual information on a scale previously unattainable, opening new avenues for a range of applications. This paper investigates the applicability and effectiveness of prompt-engineered Gemini-pro LMMs versus fine-tuned Vision Transformer (ViT) models in addressing critical security challenges. We focus on two distinct tasks: a visually evident task of detecting simple triggers, such as small squares in images, indicative of potential backdoors, and a non-visually evident task of malware classification through visual representations. Our results highlight a significant divergence in performance, with Gemini-pro falling short in accuracy and reliability when compared to fine-tuned ViT models. The ViT models, on the other hand, demonstrate exceptional accuracy, achieving near-perfect performance on both tasks. This study not only showcases the strengths and limitations of prompt-engineered LMMs in cybersecurity applications but also emphasizes the unmatched efficacy of fine-tuned ViT models for precise and dependable tasks.
keywords:
Large Multimodal Models , Vision Transformers , Fine-Tuning , Prompt Engineering , Trigger detection , Malware Classification
@@ 1 Introduction
The rise of Large Language Models (LLMs) has marked a significant milestone in the field of natural language processing, demonstrating the capability of deep learning models to understand and generate complex texts with unprecedented accuracy [1]. Building on this success, the development of Large Multimodal Models (LMMs) represents a significant leap forward, enabling the simultaneous processing of text and images [2]. Among these, Gemini-pro represents a notable attempt to leverage the vast potential of LMMs for interpreting complex content [3]. However, the application of such models in specialized domains, particularly in cybersecurity, remains in its early stages.
@@
This study aims to evaluate the efficacy of prompt-engineered LMMs, such as Gemini-pro models, compared to fine-tuned Vision Transformer (ViT) models [4] - a transformer-based architecture that deals with images- in the cybersecurity domain. The research focuses on two cybersecurity tasks: a visually evident task, which can be resolved through direct observation of images, and a non-visually evident task, which necessitates the analysis of complex visual patterns for effective outcomes.
@@
The visually evident task involves detecting specific triggers - in our case, small white squares within images - which serve as indicators of potential adversarial attacks [5]. This task is critical for the early detection of threats that could compromise the integrity of machine learning (ML) models through backdoor attacks. For this task, we used the MNIST dataset [6], augmented with a replica dataset that includes these triggers, to evaluate the model performance.
@@
The non-visually evident task involves classifying malware solely from visual representations of files, requiring the analysis of intricate visual patterns to accurately identify malware types and families based on their visual signatures. This task is paramount in enhancing the detection and mitigation of malware threats. For this task, the MaleVis dataset was utilized [7], providing a set of visual representations of malware samples divided between 25 classes and 5 families.
These tasks were selected to cover a spectrum of challenges in cybersecurity, from simple visual cues to complex visual patterns that represent real-world scenarios where visual analysis is crucial.
@@
The findings reveal a nuanced landscape of model performance across the two tasks. For trigger detection, the fine-tuned ViT models showcased remarkable proficiency, achieving 100% accuracy in identifying adversarial triggers. This performance is a testament to their ability to focus on relevant visual features. In contrast, the Gemini-pro model, despite iterative improvements through enhanced prompting, achieved a peak accuracy of 77.2%, indicating limitations in its ability to consistently recognize simple adversarial cues, even with explicit guidance through detailed prompting. For malware classification, the fine-tuned ViT models demonstrated superior performance, with accuracy rates of 97.12% for classifying among 25 malware types and 98.00% for classifying among 5 malware families. These results underscore the ViT’s robustness in discerning intricate visual patterns. Conversely, the Gemini-pro model struggled significantly, with an accuracy of 21.2% in the best-case scenario, highlighting the challenges in applying LMMs to tasks requiring deep, visual pattern analysis. These results indicate that while LMMs optimized through prompt engineering are accessible and user-friendly, their applicability in addressing specific problems is not guaranteed. This finding underscores the necessity for a critical evaluation of their use in practical scenarios. In contrast, ViT models, when fine-tuned, demonstrate superior performance, achieving notable results with a comparatively reduced parameter count.
@@
  • 1.
    Offering a unique comparative analysis between the performance of prompt-engineered LMMs and fine-tuned ViTs.
  • 2.
    Highlighting the limitations of prompt-engineered LMMs in cybersecurity applications, such as trigger detection and visual malware classification.
  • 3.
    Demonstrating the efficacy of fine-tuned ViT models across a spectrum of security tasks, including trigger detection and visual malware classification.
@@ 2 Background and Preliminaries
This section delves deeper into the foundational concepts, technologies, and tasks that underpin this comparative study of LMMs and ViT models for cybersecurity applications.
@@ 2.1 Large Multimodal Models
LMMs represent a significant advancement in integrating visual and textual data analysis, pushing the boundaries of traditional computer vision [8]. These models, exemplified by Gemini-pro, are characterized by their large-scale neural network architectures, which enable them to process and interpret both visual and textual information comprehensively. This dual capability allows LMMs to provide more contextually rich and nuanced insights than models limited to a single data type. LMMs are particularly effective in applications that require a blend of visual understanding and textual interpretation, such as in scenarios where a prompt (text) guides the model on how to analyze a provided image. In this study, we investigate the efficacy of utilizing LMMs for security applications, exploring their potential to enhance cybersecurity tasks.
2.2 Vision Transformer Models
@@ 2.3 Prompt Engineering
Prompt engineering is a technique that involves the strategic formulation of input prompts to guide the behavior of AI models, particularly those based on the transformer architecture [13]. In the context of LMMs, effective prompt engineering can significantly influence the model’s focus, interpretation, and decision-making processes. This approach is especially relevant for applications where adapting a model to specific tasks is necessary, including in cybersecurity. By carefully designing prompts that specify or hint at the desired output, practitioners can leverage the model’s pre-trained knowledge base to achieve remarkable performance on a range of tasks without having to retrain the architecture. This is why this technique is gaining popularity, as it minimizes the involvement in model training and maintenance.
@@ 2.4 Fine-Tuning
Fine-tuning is a critical process in adapting pre-trained models to new tasks or domains. It involves the adjustment of a model’s parameters on a specific dataset, allowing the model to tailor its pre-existing knowledge to the nuances of the new task [14]. This process is essential for optimizing model performance in specialized applications. Fine-tuning is particularly advantageous in situations where data for the new task is limited, as it allows for the efficient transfer of knowledge from a related but more extensive dataset. In our study, fine-tuning ViT models for cybersecurity tasks such as trigger detection and visual malware classification is explored to investigate the potential of this approach in enhancing model specificity and accuracy.
@@ 2.5 Cybersecurity Tasks
This research focuses on two cybersecurity tasks: trigger detection, a visually evident task, that can be resolved through direct observation of images, and malware classification, a non-visually evident task, that necessitates the analysis of complex visual patterns for effective outcomes.
@@ 2.5.1 Trigger Detection
Trigger detection plays a pivotal role in identifying and mitigating backdoor attacks in ML systems [15]. A backdoor in an ML model is a maliciously inserted vulnerability that causes the model to behave incorrectly when triggered by a specific input pattern while functioning normally otherwise [16]. These triggers, often inconspicuous to human observers, can be embedded within the training data as small, strategically placed artifacts or patterns. Once the model is trained on this poisoned dataset, the attacker can activate the backdoor at will by including the trigger in any input, leading to a predetermined incorrect output.
@@ 2.5.2 Malware Classification
The task of malware classification represents a critical challenge in cybersecurity, aiming to accurately identify and categorize malicious software based on its digital footprint [18]. Traditional approaches to malware classification often rely on signature-based or behavior-based analysis, requiring continuous updates to detect new malware variants [19]. However, the emergence of ML and computer vision technologies has introduced a novel approach: visual-based malware classification [20]. This method converts binary data from software files into visual representations, such as images, which can then be analyzed using image recognition techniques to identify patterns indicative of malicious code. This approach to malware classification leverages the deep learning frameworks’ capacity to recognize complex patterns in visual data. By analyzing the visual signatures of malware samples, models can learn to distinguish between various malware types and legitimate software, offering a scalable and adaptive solution for cybersecurity defenses. The use of visual analysis for malware classification is particularly promising due to its ability to generalize across different malware variants, potentially reducing the reliance on frequently updated signature databases. In our study, we employ the MaleVis dataset to evaluate the performance of fine-tuned ViT models and prompt-engineered LMMs in classifying malware based on these visual representations.
@@ 3 Related Work
LMMs represent a transformative trend in the realm of artificial intelligence and ML. These models, capable of processing and extracting insights from both text and images, have gained increasing prominence across various domains, offering a versatile approach to analyzing visual data [21, 8]. These models enable users to simply send an image and request comprehensive analysis or insights derived from it, heralding a new era of versatility and convenience [22, 13]. Notably, Google’s release of its multimodal model, Gemini, has made LMMs accessible for free via API, thus democratizing their utilization and empowering researchers and practitioners to integrate them into various applications [3]. While Gemini has demonstrated remarkable performance across numerous domains [23, 24, 25, 26], its application in security contexts remains largely unexplored. Historically, the literature has abounded with task-specific vision models tailored for security applications, including threat detection [27, 28], intrusion detection [29, 30], malware classification [31, 32], and facial recognition [33, 34].
However, with the ease of leveraging the extensive knowledge encapsulated within LMMs through simple API calls, there arises a compelling question: can LMMs, such as Gemini, effectively supplant the need for task-specific, fine-tuned ViTs in cybersecurity applications? In the realm of cybersecurity, specifically phishing detection, a recent study comparing prompt engineering to fine-tuning in LLMs has shown that while fine-tuning leads to better performance, prompt engineering still yields decent outcomes [35].
This study aims to explore the potential of LMMs, represented by Gemini, in the realm of cybersecurity, assess their performance in critical security tasks like trigger detection and malware classification, and ascertain whether they could emerge as viable alternatives to ViTs fine-tuned for specific tasks. In doing so, this research endeavors to contribute novel insights into the comparative utility of LMMs and fine-tuned ViTs, thereby informing decisions regarding their adoption in future security applications.
@@ 4 Methodology
In this section, we delineate the methodology adopted in our study to assess the performance of prompt-engineered LMMs and fine-tuned ViT models in executing two distinct cybersecurity tasks.
@@ 4.2 Fine-tuning ViT Models
Fine-tuning pre-trained models involves adjusting their weights to specialize them for a specific task. In our study, the fine-tuning of Vision Transformer (ViT) models involves the following steps:
@@ 5.1.1 Datasets
  • 1.
    Trigger Detection: For the task of trigger detection, we utilized the MNIST dataset, a standard benchmark in the field of ML for handwritten digit recognition, containing 70,000 samples. To simulate adversarial triggers, we generated a modified version of the MNIST dataset by adding a small white square (4x4 pixels) at a randomly chosen corner of each image as shown in Figure 3. This modification created a dataset that mimics the presence of subtle, adversarial manipulations aimed at compromising ML models through backdoor attacks. We combined the original and modified datasets, selecting 500 samples (equally balanced between triggered and untriggered images) as a test set. These are used to evaluate the performance of the Gemini-pro model and the fine-tuned ViT model. The actual fine-tuning of the ViT is done using the rest of the data.
@@
  • 2.
    Malware Classification: For the task of malware classification, we utilized the MaleVis dataset [7], an open-set image dataset comprising 25 malware classes that can be grouped into 5 malware families. Designed for vision-based multi-class malware recognition research, this dataset features RGB images derived from binary data of malware files. It includes a total of 12,393 images, categorized as detailed in Table 1, offering a wide range of visual representations for extensive model training and evaluation. The dataset facilitates two tasks: predicting 5 malware families and 25 individual malware classes. For each task, a test set with 50 samples per class was reserved, resulting in 250 samples for malware family prediction and 1,250 samples for malware class prediction. These test sets are utilized to assess the performance of prompt-engineered LMMs and fine-tuned ViTs. For the fine-tuning of ViTs, we employed the training portion of the dataset.
@@ 5.1.2 Models
  • 1.
    Gemini-pro: For the LMM approach, we employed the Gemini-pro model as it is the sole model accessible via the API and facilitates seamless integration with various applications. Within this framework, we executed a series of prompt-engineering strategies to direct the model’s analysis of images for both tasks, spanning from minimal guidance in baseline prompts to specific instructions provided in enhanced prompts. Further details will be elaborated upon in the subsequent sections.
  • 2.
    Fine-tuned ViT: For the ViT models, we applied fine-tuning techniques to adapt the pre-trained ViT model (’google/vit-base-patch16-224’) to our specific classification tasks. The fine-tuning process involved adjusting the model on a subset of the task-specific datasets (trigger detection and malware classification) to optimize its performance for these particular applications. The training process for the ViT model was carefully monitored using a validation set (10% of the training size) to prevent overfitting, ensuring that the model generalized well to new, unseen data.

### v2

@@ Abstract
The success of Large Language Models (LLMs) has led to a parallel rise in the development of Large Multimodal Models (LMMs), which have begun to transform a variety of applications. These sophisticated multimodal models are designed to interpret and analyze complex data by integrating multiple modalities such as text and images, thereby opening new avenues for a range of applications. This paper investigates the applicability and effectiveness of prompt-engineered LMMs that process both images and text, including models such as LLaVA, BakLLaVA, Moondream, Gemini-pro-vision, and GPT- 4o, compared to fine-tuned Vision Transformer (ViT) models in addressing critical security challenges. We focus on two distinct security tasks: 1) a visually evident task of detecting simple triggers, such as small pixel variations in images that could be exploited to access potential backdoors in the models, and 2) a visually non-evident task of malware classification through visual representations. In the visually evident task, some LMMs, such as Gemini-pro-vision and GPT-4o, have demonstrated the potential to achieve good performance with careful prompt engineering, with GPT-4o achieving the highest accuracy and F1-score of 91.9% and 91%, respectively. However, the fine-tuned ViT models exhibit perfect performance in this task due to its simplicity. For the visually non-evident task, the results highlight a significant divergence in performance, with ViT models achieving F1-scores of 97.11% in predicting 25 malware classes and 97.61% in predicting 5 malware families, whereas LMMs showed suboptimal performance despite iterative prompt improvements. This study not only showcases the strengths and lim- limitations of prompt-engineered LMMs in cybersecurity applications but also emphasizes the unmatched efficacy of fine-tuned ViT models for precise and dependable tasks.
Keywords Large Multimodal Models  $\cdot$ Vision Transformers  $\cdot$ Fine-Tuning  $\cdot$ Prompt Engineering  $\cdot$ Trigger detection  $\cdot$ Malware Classification
@@ 1 Introduction
The rise of Large Language Models (LLMs) has marked a significant milestone in the field of natural language processing, demonstrating the capability of deep learning models to understand and generate complex texts with unprecedented accuracy [1]. Building on this success, the development of Large Multimodal Models (LMMs) represents a significant leap forward, enabling the simultaneous processing of multiple modalities including text, images, video, and speech [2, 3]. Among these, Gemini-pro-vision [4], LLaVA [5], and GPT-4V [6] represent notable attempts to leverage the vast potential of LMMs for interpreting complex content based on images and text. These models have been utilized in various applications [7, 8], offering the advantage of functioning as black boxes with prompt engineering. This facilitates application development and eliminates the need for training specialized models for specific tasks. However, the application of such models in specialized domains, particularly in cybersecurity, remains largely unexplored.
@@
This study evaluates the effectiveness of prompt-engineered LMMs versus fine-tuned Vision Transformer (ViT) models [9], which are transformer-based architectures designed for processing images. The research focuses on two cybersecurity tasks for evaluation: one visually evident, which can be resolved through direct observation of images, and another visually non-evident, which requires analysis of complex visual patterns for effective outcomes.
@@
The visually evident task involves detecting specific triggers or markers, notably small pixel variations in images. These variations aim to exploit potential backdoors in machine learning (ML) models [10]. Early detection of these variations is crucial for maintaining the integrity of ML models. For this task, we used the MNIST dataset [11], augmented with a replica dataset that includes these triggers, to evaluate the model performance.
@@
The visually non-evident task involves classifying malware solely from visual representations of files, requiring the analysis of intricate visual patterns to accurately identify malware types and families based on their visual signatures. This task is paramount in enhancing the detection and mitigation of malware threats. For this task, the MaleVis dataset was utilized [12], providing a set of visual representations of malware samples divided between 25 classes and 5 families.
@@
The findings reveal a nuanced landscape of model performance across the two tasks. For trigger detection, the fine-tuned ViT showcased exceptional proficiency, achieving 100% accuracy in identifying adversarial triggers. This performance is a testament to their ability to focus precisely on relevant visual features. In contrast, LMMs, through iterative improvements with enhanced prompting, achieved a peak accuracy of 91.9% and an F1-score of 91% with GPT-4o, indicating a solid performance that, while effective, does not reach the exceptional standard set by the fine-tuned ViT models. For malware classification, the fine-tuned ViT demonstrated superior performance, with F1-scores of 97.11% for classifying among 25 malware types and 97.61% for classifying among 5 malware families. These results underscore the ViT’s robustness in discerning intricate visual patterns. Conversely, LMMs faced challenges in this more demanding task, where Gemini-pro-vision achieved an F1-score of 12.7%, and GPT-4o reached an F1-score of 34.6% in the best-case scenario. This highlights the difficulties LMMs encounter in tasks requiring deep visual pattern analysis. These results suggest that while LMMs optimized through prompt engineering are accessible and user-friendly, their applicability in addressing specific problems is not guaranteed. This finding underscores the necessity for a critical evaluation of their use in practical scenarios. In contrast, ViT models, when fine-tuned, demonstrate superior performance, achieving notable results with a comparatively reduced parameter count.
@@
  • Introducing the application of LMMs to cybersecurity applications, marking the first application of these models in this specific context.
  • Comparing the performance of prompt-engineered LMMs with fine-tuned ViT models.
  • Identifying the limitations of prompt-engineered LMMs in cybersecurity tasks such as trigger detection and visual malware classification.
  • Demonstrating the effectiveness of fine-tuned ViT models across image-based security tasks.
@@ 2 Background and Preliminaries
This section examines the details of foundational concepts, technologies, and tasks that underpin this comparative study of prompt-engineered LMMs and ViT models for cybersecurity applications.
@@ 2.1 Large Multimodal Models
LMMs represent a significant advancement in integrating visual and textual data analysis, pushing the boundaries of traditional computer vision [8]. These models are characterized by their large-scale neural network architectures, which enable them to process and interpret both visual and textual information comprehensively. This dual capability allows LMMs to provide more contextually rich and nuanced insights than models limited to a single data type. LMMs are particularly effective in applications that require a blend of visual understanding and textual interpretation, such as in scenarios where a prompt (text) guides the model to analyze a provided image. In this study, we investigate the efficacy of utilizing LMMs for security applications, exploring their potential to enhance cybersecurity tasks.
2.2 Vision Transformer (ViT)
@@ 2.3 Prompt Engineering
Prompt engineering is a technique that involves the strategic formulation of input prompts to guide the behavior of AI models, particularly those based on the transformer architecture [17]. In the context of LMMs, effective prompt engineering can significantly influence the model’s focus, interpretation, and decision-making processes. This approach is especially relevant for applications where adapting a model to specific tasks is necessary. By carefully designing prompts that specify or hint at the desired output, practitioners can leverage the model’s pre-trained knowledge base to achieve remarkable performance on a range of tasks without having to retrain the architecture. This is why this technique is gaining popularity, as it minimizes the involvement in model training and maintenance.
@@ 2.4 Fine-Tuning
Fine-tuning is a critical process in adapting pre-trained models to new tasks or domains. It involves adjusting a model’s parameters on a specific dataset, enabling the model to tailor its pre-existing knowledge to the nuances of the new task [18]. This process is essential for optimizing model performance in specialized applications. Fine-tuning is particularly advantageous in situations where data for the new task is limited, as it allows for the efficient transfer of knowledge from a related but larger dataset. In this study, we explore the fine-tuning of ViT models for cybersecurity tasks such as trigger detection and visual malware classification, to investigate their potential to achieve high performance compared to prompt-engineered LMMs.
@@ 2.5 Cybersecurity Tasks
This research focuses on two cybersecurity tasks: trigger detection, a visually evident task, that can be resolved through direct observation of images, and malware classification, a visually non-evident task, that necessitates the analysis of complex visual patterns for effective outcomes.
@@ 2.5.1 Trigger Detection
Trigger detection plays a pivotal role in identifying and mitigating backdoor attacks in ML systems [19]. A backdoor in an ML model is a maliciously inserted vulnerability that causes the model to behave incorrectly when triggered by a specific input pattern while functioning normally otherwise [20]. These triggers, often inconspicuous to human observers, can be embedded within the training data as small, strategically placed artifacts or patterns. Once the model is trained on this poisoned dataset, the attacker can activate the backdoor at will by including the trigger into any input, leading to a predetermined incorrect output.
@@ 2.5.2 Malware Classification
The task of malware classification represents a critical challenge in cybersecurity, aiming to accurately identify and categorize malicious software based on its digital footprint [22]. Traditional approaches to malware classification often rely on signature-based or behavior-based analysis, requiring continuous updates to detect new malware variants [23]. However, the emergence of ML and computer vision technologies has introduced a novel approach: visual-based malware classification [24]. This method converts binary data from software files into visual representations, such as images, which can be analyzed using image recognition techniques to identify patterns indicative of malicious code. This approach to malware classification leverages the deep learning frameworks’ capacity to recognize complex patterns in visual data. By analyzing the visual signatures of malware samples, models can learn to distinguish between various malware types and legitimate software, offering a scalable and adaptive solution for cybersecurity defenses. The use of visual analysis for malware classification is particularly promising due to its ability to generalize across different malware variants, potentially reducing the reliance on frequently updated signature databases. In this study, we employ the MaleVis dataset to evaluate the performance of fine-tuned ViT models and prompt-engineered LMMs in classifying malware based on these visual representations.
@@ 3 Related Work
LMMs represent a transformative trend in artificial intelligence and machine learning. These models, capable of processing and extracting insights from both text and images, have gained prominence across various domains, offering a versatile approach to analyzing visual data [25, 8]. LMMs enable users to upload an image and request comprehensive analysis or insights derived from it, introducing a new dimension of versatility and convenience [26, 17]. Numerous LMMs exist that can process both visual and textual information and can be utilized via simple API calls, simplifying their use and empowering researchers and practitioners to integrate them into diverse applications. Notable examples include LLaVA [5], Gemini-pro-vision [4], and GPT-4o [27]. While these models have demonstrated remarkable performance across many domains [28, 29, 30, 31, 32, 33, 34, 35], their application in security contexts remains largely unexplored. Recent studies have leveraged LLMs for various cybersecurity tasks [36, 37, 38, 39, 40, 41], yet similar explorations with LMMs do not exist up to our knowledge, making this study the first to venture into this field. Historically, the literature has abounded with task-specific vision models tailored for security applications, such as threat detection [42, 43], intrusion detection [44, 45], malware classification [46, 47], and facial recognition [48, 49]. However, with the ease of leveraging the extensive knowledge encapsulated within LMMs through simple API calls, there arises a compelling question: Can LMMs effectively replace the need for task-specific, fine-tuned Vision Transformers (ViTs) in cybersecurity applications? ViTs, with their underlying transformer architecture, share similarities with LMMs, prompting an investigation into their comparative performance. This study aims to explore the potential of LMMs in the realm of cybersecurity, assess their performance in critical security tasks like trigger detection and malware classification, and ascertain whether they could emerge as viable alternatives to fine-tuned ViTs. In doing so, this research attempts to contribute novel insights into the comparative utility of LMMs and ViTs, thereby informing decisions regarding their adoption in future security applications.
@@ 4 Methodology
In this section, we delineate the methodology adopted to assess the performance of prompt-engineered LMMs and fine-tuned ViT models in executing two distinct cybersecurity tasks.
@@ 4.2 Fine-tuning ViT Models
Fine-tuning pre-trained models involves adjusting their weights to specialize them for a specific task. In this study, the fine-tuning of ViT models involves the following steps:
@@ 5.1.1 Datasets
  • Trigger Detection: For the task of trigger detection, we utilized the MNIST dataset, a standard benchmark in the field of ML for handwritten digit recognition, containing 70,000 samples. To simulate adversarial triggers, we generated a modified version of the MNIST dataset by adding a small white square (4x4 pixels) at a randomly chosen corner of each image as shown in Figure 3. This modification created a dataset that mimics the presence of subtle, adversarial pixel manipulations aimed at compromising ML models through backdoor attacks. We combined the original and modified datasets, selecting 500 samples (equally balanced between triggered and untriggered images) as a test set. These are used to evaluate the performance of the various LMMs and the fine-tuned ViT model. The actual fine-tuning of the ViT is done using the rest of the data.
@@
  • Malware Classification: For the task of malware classification, we utilized the MaleVis dataset [12], an image dataset comprising 25 malware classes that can be grouped into 5 malware families. Designed for vision-based multi-class malware recognition research, this dataset features RGB images derived from binary data of malware files. It includes a total of 12,393 images, categorized as detailed in Table 1, offering a wide range of visual representations for extensive model training and evaluation. The dataset facilitates two tasks: predicting 5 malware families and 25 individual malware classes. For each task, a test set with 50 samples per class was reserved, resulting in 250 samples for malware family prediction and 1,250 samples for malware class prediction. These test sets are utilized to assess the performance of prompt-engineered LMMs and fine-tuned ViTs. For the fine-tuning of ViTs, we employed the training portion of the dataset.
@@ 5.1.2 Models
  • LMMs: For prompt-engineered LMMs, we employed 5 distinct models including: LLaVA, BakLLaVA, Moondream2, Gemini-pro-vision, and GPT-4o. The first three models are accessible via Hugging Face and the last two are accessible via API, which facilitates seamless integration with various applications. Within this framework, we executed a series of prompt-engineering strategies to direct the models’ analysis of images for both tasks, spanning from minimal guidance in baseline prompts to specific instructions provided in enhanced prompts. Further details will be elaborated upon in the subsequent sections.
  • Fine-tuned ViT: For the ViT models, we applied fine-tuning techniques to adapt the pre-trained ViT model (’google/vit-base-patch16-224’) to our specific classification tasks. The fine-tuning process involved adjusting the model on a subset of the task-specific datasets (trigger detection and malware classification) to optimize its performance for these particular applications. The training process for the ViT model was carefully monitored using a validation set (10% of the training size) to prevent overfitting, ensuring that the model generalizes well to new, unseen data.